OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). x=4y 3x-2y=20

OpenStudy (anonymous):

3x-2y=20 3(4y)-2y=20 12y-2y=20 10y=20 y =20/10 y =2

OpenStudy (anonymous):

3x-2y=20 3x-2(2)=20 3x-4=20 3x=20+4 x=24/3 x=8